Suppose in the country of Lioness, the following working-age demographic data are available.

Unemployment 10,000

Labor force 150,000

Not in the labor force 60,000

a) How many people are employed in Lioness?

b) What is the unemployment rate?

c) What is the participation rate?
Answer #1

ANSWER:

A) No of people employed = labor force - unemployment = 150,000 - 10,000 = 140,000

B) Unemployment rate = (unemployment / labor force) * 100 = (10,000 / 150,000) * 100 = 0.667 * 100 = 6.67%

C) Participation rate = (labor force / population) * 100

population = labor force + not in labor force = 150,000 + 60,000 = 210,000

Participation rate = (150,000 / 210,000) * 100 = 0.7142 * 100 = 71.42%